A local pizzeria sells 500 large pepperoni pizzas per week at a price of $20 each. Suppose the owner of the pizzeria tells you that the price elasticity of demand for his pizza is -2, and he asks you for advice. He wants to know two things. First, how many pizzas will he sell if he cuts his price by 10%? Second, how will his revenue be affected?
1
Expert's answer
2017-10-11T12:22:06-0400
a) E = (Q1-Q/Q)/(P1-P)/P, Q1-Q/Q = -2*(18-20)/20 = 0,2 => Q1 = 600.
b) I = Q1 * P1 = 600 * 18 =1080. Then total revenue will be more than previous price (I =Q0 * P0 = 1000).